# Bitcoin Cross-Chain Bridge: A Taxonomy and Its Promise in Artificial Intelligence of Things

### v1

@@ Abstract
Bitcoin’s limited scripting capabilities and lack of native interoperability mechanisms have constrained its integration into the broader blockchain ecosystem, especially decentralized finance (DeFi) and multi-chain applications. This paper presents a comprehensive taxonomy of Bitcoin cross-chain bridge protocols, systematically analyzing their trust assumptions, performance characteristics, and applicability to the Artificial Intelligence of Things (AIoT) scenarios. We categorize bridge designs into three main types: naive token swapping, pegged-asset bridges, and arbitrary-message bridges. Each category is evaluated across key metrics such as trust model, latency, capital efficiency, and DeFi composability. Emerging innovations like BitVM and recursive sidechains are highlighted for their potential to enable secure, scalable, and programmable Bitcoin interoperability. Furthermore, we explore practical use cases of cross-chain bridges in AIoT applications, including decentralized energy trading, healthcare data integration, and supply chain automation. This taxonomy provides a foundational framework for researchers and practitioners seeking to design secure and efficient cross-chain infrastructures in AIoT systems.
@@ 1 Introduction
Blockchain technology, as a decentralized ledger, has been widely used in various decentralized AIoT applications, such as smart healthcare [21, 22, 23], supply chains [24, 25, 26], and energy trading scenarios [27, 28, 29]. Depending on the specific entities and solutions involved, various blockchain ledgers may be implemented. Ensuring interoperability among these heterogeneous ledgers presents a considerable challenge to AIoT applications. The cross-chain bridges technology provides a feasible connection among those distributed ledgers.
@@ 2 Background
2.1 Overview of Cross-Chain Bridges
@@
2.2 Naive Tokens Swapping
@@
2.3 Pegged-Asset Bridges
@@
2.4 Message-Arbitrary Bridges
@@
{forest}
[…]
@@ 3 Literature Review
3.1 Atomic Swaps and HTLCs
@@ 3.7 Sidechains
3.8 Trust Model of Bridges
@@
4 Bridging the Cross Blockchains in AIoT Applications
@@
4.1 Case in Energy Trading
@@ 4.2 Case in Healthcare
4.3 Case in Supply Chain
Global supply chains often involve multiple platforms, such as a shipping AIoT platform, a trade finance blockchain, or a customs paperwork ledger [24, 26]. The neutrality and liquidity of cryptocurrencies, such as ETH, make them attractive as a settlement currency to bridge these systems. Collateralized vaults enable supply chain participants to utilize ETH for trade transactions across networks. For example, an exporter could lock its ETH within a vault to mint a token on a supply-chain-specific blockchain that monitors shipments. This token can be used within that network’s smart contracts—potentially serving as escrow within an AIoT-supervised letter-of-credit contract. Upon the shipment’s arrival and subsequent confirmation of delivery by AIoT oracles, the escrow is released to the exporter. Due to the token being pegged, the exporter can later redeem it for actual ETH, thereby obtaining global liquidity. Meanwhile, the importer is able to transact in ETH without directly interacting with the Ethereum main chain during the procedure. Collateralized pegs give supply chain platforms a decentralized and robust common currency, resolving interoperability challenges between financial and physical flows. However, the trade-off, as noted in Section 3.3, involves the management of vault infrastructure and the requirement for over-collateralization, which introduces added complexity and expense. Participants must evaluate these factors against the advantage of avoiding dependence on a single central escrow or foreign exchange intermediary for cross-network transactions.
5 Conclusion

### v2

@@ Abstract
Bitcoin’s limited programmability constrains its integration with DeFi and multi-chain applications. This paper presents a comprehensive taxonomy of Bitcoin cross-chain bridge protocols, systematically analyzing their trust assumptions, performance characteristics, and applicability to the Artificial Intelligence of Things (AIoT) scenarios. We categorize bridge designs into three main types: naive token swapping, pegged-asset bridges, and arbitrary-message bridges. Each category is evaluated across key metrics such as trust model, latency, capital efficiency, and DeFi composability. Emerging innovations like BitVM and recursive sidechains are highlighted for their potential to enable secure, scalable, and programmable Bitcoin interoperability. Furthermore, we explore practical use cases of cross-chain bridges in AIoT applications, including decentralized energy trading, healthcare data integration, and supply chain automation. This taxonomy provides a foundational framework for researchers and practitioners seeking to design secure and efficient cross-chain infrastructures in AIoT systems.
@@ I Introduction
Blockchain technology, as a decentralized ledger, has been widely used in various decentralized AIoT applications, such as smart healthcare [21, 22, 23], supply chains [24, 25, 26, 27], and energy trading scenarios [28, 29, 30]. Depending on the specific entities and solutions involved, various blockchain ledgers may be implemented. Ensuring interoperability among these heterogeneous ledgers presents a considerable challenge to AIoT applications. The cross-chain bridges technology provides a feasible connection among those distributed ledgers.
@@ II Background
II-A Overview of Cross-Chain Bridges
@@
II-B Naive Tokens Swapping
@@
II-C Pegged-Asset Bridges
@@
II-D Message-Arbitrary Bridges
@@
Cross-Chain Bridges Naive Token Swapping Atomic Swaps (HTLC) Payment Channel Pegged-Asset Bridges Federated Sidechains Collateralized Vault BitVM Arbitrary-Message Bridges On-chain Relays External Verifier
@@ III Literature Review
III-A Atomic Swaps and HTLCs
@@ III-G Sidechains
III-H Trust Model of Bridges
@@
IV Bridging the Cross Blockchains in AIoT Applications
@@
IV-A Case in Energy Trading
@@ IV-B Case in Healthcare
IV-C Case in Supply Chain
Global supply chains often involve multiple platforms, such as a shipping AIoT platform, a trade finance blockchain, or a customs paperwork ledger [24, 25]. The neutrality and liquidity of cryptocurrencies, such as ETH, make them attractive as a settlement currency to bridge these systems. Collateralized vaults enable supply chain participants to utilize ETH for trade transactions across networks. For example, an exporter could lock its ETH within a vault to mint a token on a supply-chain-specific blockchain that monitors shipments. This token can be used within that network’s smart contracts—potentially serving as escrow within an AIoT-supervised letter-of-credit contract. Upon the shipment’s arrival and subsequent confirmation of delivery by AIoT oracles, the escrow is released to the exporter. Due to the token being pegged, the exporter can later redeem it for actual ETH, thereby obtaining global liquidity. Meanwhile, the importer is able to transact in ETH without directly interacting with the Ethereum main chain during the procedure. Collateralized pegs give supply chain platforms a decentralized and robust common currency, resolving interoperability challenges between financial and physical flows. However, the trade-off, as noted in Section III-C, involves the management of vault infrastructure and the requirement for over-collateralization, which introduces added complexity and expense. Participants must evaluate these factors against the advantage of avoiding dependence on a single central escrow or foreign exchange intermediary for cross-network transactions.
V Conclusion